Considering the 5,800 enrollees in the College of Education, a researcher would like to make a study with 97.25% precision. How many samples must be taken?
1
Expert's answer
2020-11-18T18:50:00-0500

We build 97.25% confidence interval for population mean or population proportion. This confidence interval covers population mean or population proportion with the probability 97.25%.

If we build 97.25% confidence interval for population mean:

Lower limit = "M-z_{1.375}\sigma_M"

Upper limit = "M+z_{1.375}\sigma_M"

Margin of error "ME=z_{1.375}\sigma_M"

"\sigma_M=\frac{\sigma}{\sqrt{N}}"

"M\text{ --- sample mean}\\
\sigma\text{ --- population standard deviation}\\
N\text{ --- sample size}\\
z_{1.375}\approx 0.0838\text{ --- the corresponding critical value}"

Then we have:

"N=\frac{z_{1.375}^2\sigma^2}{ME^2}"

If we build 97.25% confidence interval for population proportion:

"N=\frac{z_{1.375}^2p(1-p)}{ME^2}"

"p\text{ --- estimated proportion of the population which}\\
\text{has the attribute in question}."